# Relationship between the ray variability and the pc-scale jet in the blazar 3C 454.3

### v1

@@ Acknowledgements.
EP would like to express her sincere gratitude to the faculty and graduate students of the Max Planck Institute for Radio Astronomy (MPIfR) for the insightful discussions and generous hospitality extended during her visit. Particular thanks are extended to Dr. D. Kim for his dedication of time and for engaging in insightful discussions concerning the analysis of ray data. We thank Dr. G. Paraschos for his role as the internal referee at the MPIfR. This work is supported by CONAHCyT research grants 280789 and 320987, and by the MPIfR-Mexico Max Planck Partner Group led by VMP-A. This publication is part of the M2FINDERS project which has received funding from the European Research Council (ERC) under the European Unionās Horizon 2020 Research and Innovation Programme (grant agreement No. 101018682). This research has made use of data from the OVRO 40 m monitoring program (Richards etĀ al., 2011), which is supported in part by NASA grants NNX08AW31G, NNX11A043G, and NNX14AQ89G and NSF grants AST-0808050 and AST-1109911. This research has used data from the MOJAVE database, which is maintained by the MOJAVE team (Lister etĀ al., 2009, 2018). This study makes use of VLBA data from the BEAM-ME and VLBA-BU-BLAZAR Monitoring Programs ( http://www.bu.edu/blazars/BEAM-ME.html), funded by NASA through the Fermi Guest Investigator Program. The VLBA is an instrument of the National Radio Astronomy Observatory. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated by Associated Universities, Inc. The Submillimeter Array is a joint project between the Smithsonian Astrophysical Observatory and the Academia Sinica Institute of Astronomy and Astrophysics and is funded by the Smithsonian Institution and the Academia Sinica. We recognize that Maunakea is a culturally important site for the indigenous Hawaiian people; we are privileged to study the cosmos from its summit. Software. Astropy (Astropy Collaboration etĀ al., 2013, 2018), CASA (McMullin etĀ al., 2007), Fermitools v 2.0.8 (https://fermi.gsfc.nasa.gov/ssc/data/analysis/).References

### v2

@@ Acknowledgements.
EP would like to express her sincere gratitude to the faculty and graduate students of the Max Planck Institute for Radio Astronomy (MPIfR) for the insightful discussions and generous hospitality extended during her visit. Particular thanks are extended to Dr. D. Kim for his dedication of time and for engaging in insightful discussions concerning the analysis of -ray data. We thank Dr. G. Paraschos for his role as the internal referee at the MPIfR. This work is supported by CONAHCyT research grants 280789 and 320987, and by the MPIfR-Mexico Max Planck Partner Group led by VMP-A. This publication is part of the M2FINDERS project which has received funding from the European Research Council (ERC) under the European Unionās Horizon 2020 Research and Innovation Programme (grant agreement No. 101018682). This research has made use of data from the OVRO 40 m monitoring program (Richards etĀ al., 2011), which is supported in part by NASA grants NNX08AW31G, NNX11A043G, and NNX14AQ89G and NSF grants AST-0808050 and AST-1109911. This research has used data from the MOJAVE database, which is maintained by the MOJAVE team (Lister etĀ al., 2009, 2018). This study makes use of VLBA data from the BEAM-ME and VLBA-BU-BLAZAR Monitoring Programs ( http://www.bu.edu/blazars/BEAM-ME.html), funded by NASA through the Fermi Guest Investigator Program. The VLBA is an instrument of the National Radio Astronomy Observatory. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated by Associated Universities, Inc. The Submillimeter Array is a joint project between the Smithsonian Astrophysical Observatory and the Academia Sinica Institute of Astronomy and Astrophysics and is funded by the Smithsonian Institution and the Academia Sinica. We recognize that Maunakea is a culturally important site for the indigenous Hawaiian people; we are privileged to study the cosmos from its summit. Software. Astropy (Astropy Collaboration etĀ al., 2013, 2018), CASA (McMullin etĀ al., 2007), Fermitools v 2.0.8 (https://fermi.gsfc.nasa.gov/ssc/data/analysis/).References